OpenStudy (anonymous):

give an example, while completing the square, where the outcome will be a negative number

OpenStudy (anonymous):

is it possible?

OpenStudy (lalaly):

\[x^2 + 6x + 10 = 0\]\[(x+3)^2-9+10=0\]\[(x+3)^2=-1\] u'll get complex solutions